OpenStudy (anonymous):

What is the solution to the rational equation (x/x^2-9) - (1/x-3) = (1/4x-12)

OpenStudy (anonymous):

you know you can write x^2 - 9 (x-3)(x+3) and 4x - 12 4(x-3) now you know how to solve it ?

OpenStudy (anonymous):

i think... would -9/7 be the right answer?

OpenStudy (anonymous):

i dont think lets do more steps : so we said : x/((x-3)(x+3)) - 1/(x-3) = 1/(4(x-3)) now we can get rid of (x-3) by saying x is not 3 so we left with : x/(x+3) - 1 = 1/4

OpenStudy (anonymous):

now multiply the equation by x+3 and solve for x

OpenStudy (anonymous):

i got -15

OpenStudy (anonymous):

its good .. you can plug it into the equation that you stated with and check if it works!

OpenStudy (anonymous):

that's something good about equations .. you can go back to the first equation and check your answer.

OpenStudy (anonymous):

it does but im not sure i did the math right, you got -15 too?

OpenStudy (anonymous):

yes its fine.

OpenStudy (anonymous):

okay perfect thanks !!

OpenStudy (anonymous):

yw :)